OpenStudy (anonymous):

Find the slope of every line that is perpendicular to the graph of the equation - y+2x=6

OpenStudy (anonymous):

Slopes of perpendicular lines multiply to give -1.

OpenStudy (bahrom7893):

perpendicular means slope is negative inverse.. calculate the slope and find it's negative inverse.

OpenStudy (anonymous):

so replace -1 with x?

OpenStudy (bahrom7893):

Okay look: -y +2x = 6 -y = -2x+6 y = 2x-6 Slope of ur original equation is 2

OpenStudy (anonymous):

In other words, slope of line * slope of perpendicular line(s) = -1

OpenStudy (bahrom7893):

Now to get the slope of perpendicular line do: -1/2

OpenStudy (amistre64):

ax + by = c is the format that you are looking at here; where the slope of this line is -a/b a perpendicular line to this one will have a slope of b/a

OpenStudy (anonymous):

i dotn get it :(!

OpenStudy (anonymous):

OK, which bit?

OpenStudy (anonymous):

everything :( i need help step by step.

OpenStudy (bahrom7893):

Okay look: -y +2x = 6 -y = -2x+6 y = 2x-6 Slope of ur original equation is 2

OpenStudy (bahrom7893):

So M=2 to get the perpendicular slope do -1/M = -1/2

OpenStudy (anonymous):

-y + 2x = 6 is a line. If we put this equation in a standard way, we can just "read" the slope from it.

OpenStudy (anonymous):

ok but is asking to find the perpendicular graph of the equation using y+2x=6 :(

OpenStudy (anonymous):

y or -y?

OpenStudy (anonymous):

Your question says -y?

OpenStudy (bahrom7893):

y or -y? good question.. please post the correct question.

OpenStudy (anonymous):

its y! lol look.

OpenStudy (bahrom7893):

yea well ur original question said -y..

OpenStudy (anonymous):

OK, when you put the questions, don't use a dash, people will think it is a - sign.

OpenStudy (bahrom7893):

Okay look: y +2x = 6 y = -2x+6 Slope of ur original equation is -2

OpenStudy (bahrom7893):

Slope of perpendicular line is (-1)/(-2) = 1/2

OpenStudy (anonymous):

Go a little bit slower, bahrom..

OpenStudy (anonymous):

im sorry :( dont yell at me. so the answer is 1/2 ?

OpenStudy (bahrom7893):

yes and estudier how much slower can i go? that;s already like every step.

OpenStudy (amistre64):

you have: a x +b y; slope = -a/b +2x +1 y slope = -2/1 perp line has slope of b/a: 1/2

OpenStudy (anonymous):

thanks dolllll <3

OpenStudy (anonymous):

scilaaaax3, do you understand how to read the slope from an equation now?